Sony Corp. purchased a new machine on Jan 1, 2014. The cost of this machine was $500,000. The company estimated that the machine would have a salvage value of $20,000 at the end of its service life. The estimated service life is 4 years and its estimated total working hours are 25,000 hours. Year-end is December 31. 1. Assuming 5,000 hours used in 2014, compute the depreciation expense for the year of 2014 under Activity method. 2. Compute the depreciation expense for the year of 2014 under Straight-line depreciation method. 3. Compute the depreciation expense for the year of 2014 under Sum-of-the-years’-digits method. 4. Compute the depreciation expense for the year of 2014 under Double-declining balance method (i.e., m=2). 5. Compute the depreciation expense for the year of 2017 under Double-declining balance method (i.e., m=2).
Answer #1

Answer -

As per given information,

Cost of machine = $500000

Salvage value = $20000

Depreciable cost = $500000 - $20000 = $480000

Estimated useful life = 4 years

Estimated total working hours = 25000 hours

1. Answer -

Assuming 5000 hours used in 2014

Using Activity method:

Depreciation expense = (Depreciable cost / Total working hours) * Hours used during the year

Depreciation expense = ($480000 / 25000) * 5000

Depreciation expense = $96000

2. Answer -

Using Straight-line depreciation method:

Depreciation expense = (Depreciable cost / Useful life)

Depreciation expense = ($480000 / 4)

Depreciation expense = $120000

3. Answer -

Using Sum of the years’ digits method:

Here,

Sum of the years' digits =4 + 3 + 2 + 1 = 10

Year 1 Year 2 Year 3 Year 4
Undepreciated useful life (years) 4 3 2 1

So,

Depreciation expense = Depreciable cost * (4 / 10)

Depreciation expense = $480000 * (4 / 10)

Depreciation expense = $192000

4 and 5. Answer -

Using Double-declining balance method:

Here.

Depreciation rate = (1 / useful life) *100 = (1 / 4) * 100 = 25%

Now,

Double-declining balance method = 2 * Cost of the asset * Depreciation rate

For year 2014

Double-declining balance method = 2 * $500000 * 25%

Double-declining balance method = $250000

For year 2015

Double-declining balance method = 2 * $250000 * 25%

Double-declining balance method = $125000

For year 2016

Double-declining balance method = 2 * $125000 * 25%

Double-declining balance method = $62500

For year 2017

Double-declining balance method = 2 * $62500 * 25%

Double-declining balance method = $31250

Year Book Value (Beginning Year) Depreciation expense Book Value (End Year)
2014 $500000 $250000 $250000
2015 $250000 $125000 $125000
2016 $125000 $62500 $62500
2017 $62500 $31250 $31250